# thanks: ISTI Fellow

### v1

@@ II.3 The likelihood of ejecta parameters for AT2017gfo given electromagnetic observations
The AT2017gfo data is originally presented in [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 19, 73, 14, 74, 16, 17]. Each sample $\vec{x}_{\theta}$ is evaluated by the MLP to produce a light-curve prediction $\hat{y}$ for every one of the grizyJHK broadband filters. We calculate the residual between the MLP prediction $\hat{y}$ and the AT2017gfo observed data in time $i$ for every band $B$ by way of the reduced-$\chi^{2}$ statistic
@@ II.4 The likelihood of source parameters for GW710817 given gravitational wave observations
Following previous studies that constrain properties of GW170817 and the equation of state from GW observations [75, 76, 77, 78], we use the RIFT parameter inference engine [79, 80, 81] to explore and evaluate the marginal likelihood $L(x)$ for a wide range of different compact binary source parameters $x$. Following the RIFT paradigm, the parameters $x$ include all intrinsic (detector-frame) quantities needed to phenomenologically characterize the binary’s inspiral: its two component masses, spins, and tidal deformabilities, assumed independent of any specific EOS model. We then use standard nonparametric interpolation techniques (here, random forests from sklearn) to interpolate the marginal likelihood to provide a continuous estimate $\hat{L}(x)$.
To be more concrete, we will employ precisely the same analysis framework and settings as used in Ref. [82]. As a brief review to establish notation, gravitational wave observations of GW170817 constrain the binary masses $m_{i}$ and tidal deformabilities $\Lambda_{i}$ for each component $i$ of the binary [1, 75, 26, 83]. The presence of matter impacts the binary’s inspiral at leading order through each component’s dimensionless tidal deformability parameter $\Lambda_{i}=(2/3)k_{2}(c^{2}r_{i}/Gm_{i})^{5}$ [84], where $k_{2}$ is the $l=2$ Love number, $m_{i}$ is the neutron star mass, and $r_{i}$ is the neutron star radius. The leading order contribution to the phase evolution of a GW inspiral is given by the weighted combination of $\Lambda_{i}$ terms
@@
The impacts of tidal effects on outgoing radiation are included in many conventional approximate phenomenological estimates of outgoing radiation from merging compact binaries. In this work, we begin with precisely the marginal likelihood data accumulated in Ref. [82]. That investigation was performed with a contemporary state-of-the-art model, IMRPhenomPv2_NRTidalv2 [85, 86], which incorporates precession physics but omits higher-order modes. This analysis was performed with open GW data for GW170817 available from GWOSC [87], using the same power spectral densities provided with GWTC-1 [88, 89], over a frequency range from $23{\rm Hz}$ to $1700{\rm Hz}$, with known sky location and source luminosity distance derived from the electromagnetically-identified host galaxy. For exploration purposes, most prior assumptions were customary (e.g., uniform in detector-frame component masses), noting that the previous investigation employed multiple spin priors and mass ratio priors to accumulate marginal likelihood information over a wide range of source parameters. Although this analysis included prior information about the alignment between the binary’s angular momentum direction and the line of sight, inferred from late-time radio afterglow observation [90, 91], this information should not significantly impact intrinsic parameters.
Because the original marginal likelihood grid accumulated in Ref. [82] did not extend to cover all of the extreme BNS configurations needed in our study, we replicated and extended this analysis to ensure that marginal likelihood data was present whenever required for joint EM/GW analysis. For simplicity, we employ the same GW model throughout for all $\tilde{\Lambda}$, keeping in mind that this model was not calibrated over the full range of $\Lambda$ that are needed to carry out our investigation.
@@ II.5 Bayesian inference
We use the RIFT framework [92] to adaptively perform Monte Carlo integration and generate samples. As previously, we employ an adaptive volume Monte Carlo integrator, following closely the approach outlined in Ref. [93]. The adaptive volume integrator allows for more efficient sampling given the higher-dimensional space explored in this work. As our inference is performed via adaptive Monte Carlo integration, the reliability of our posterior can be expressed in terms of a number of effective samples $n_{\rm eff}$. Several different conventions exist for this number; see the appendix of Ref. [81] for more detailed discussion. For this study, we terminate our analyses when $n_{\rm eff}\simeq 10^{3}$, indicating sufficient convergence.
Table 1 shows the sampling parameters used during inference, along with their associated lower and upper limits and priors. The combination of the chirp mass ${\cal M}_{c}$ and mass asymmetry $\delta$ priors provide an equivalent locally uniform prior for the binary masses $m_{1}$ and $m_{2}$. In addition to the masses, we also sample for the spins of the neutron stars $s_{\rm{1z}}$ and $s_{\rm{2z}}$. We assume that both of the neutron stars have the same radius given a 1.4 $M_{\odot}$ neutron star $R_{1.4}$, treating this radius as a proxy for the nuclear EOS. As described in Sec. II.1, we introduce parameters $\alpha_{\rm{dyn}}$ and $f_{\rm{disk}}$ that encapsulate the uncertainty in our ejected dynamical and wind mass, respectively. These parameters are sampled in addition to the light-curve systematics $\sigma_{\rm{sys}}$, effectively allowing for three separate measures of systematic uncertainty in our analysis. Finally, we sample for the viewing angle $\theta$ as in prior work (e.g. Ref. [52]). All parameters in Table 1 are assumed to be dimensionless unless their names are followed by units in brackets.
@@
$\mathcal{M}_{c}$ [$M_{\odot}$] [1.1855, 1.1865] Jointly uniform $m_{1},m_{2}$
$\delta$ ${[0,0.5]}$ Jointly uniform $m_{1},m_{2}$
@@
$R_{1.4}$ [km] [$8,25$] Uniform
$\log\alpha_{\rm{dyn}}$ [$-5,2$] Uniform
$\log f_{\rm{disk}}$ [$-5,0$] Uniform
$\beta_{\phi}$ [$-2,2$] Uniform
@@ III Results
III.1 All systematics included ($\alpha_{\rm{dyn}},f_{\rm{disk}},\beta_{\phi}$)
@@
We verify that our ejecta fit posteriors produce consistent kilonova ejecta properties in Figure 3. For all the analyses in this section, we assume that both the systematic uncertainty on the ejecta velocity and the fraction of the disk $f_{\rm{disk}}$ representing the amount of ejected wind mass are treated as free parameters. Light-curve systematic uncertainty $\sigma_{\rm{sys}}$ is treated as a free parameter for all analyses in this manuscript. We begin by examining the posteriors for the binary’s ejecta properties, primarily constrained by our EM likelihood $\mathcal{L}_{\rm{EM}}$. Figure 3 shows the predicted ejected masses, bulk velocity (attributed to the dynamical ejecta), and the viewing angle derived from the inferences presented in Figure 5. In this analysis, we adopt $v_{w}=0.10c$, consistent with numerical simulations and our previous analysis of AT2017gfo [52]. The dashed black, red, and blue lines correspond to inferences using only AT2017gfo data $\mathcal{L}_{\rm{EM}}$ for the KruFo20  DiCo20  and Nedora21 ejecta fits, respectively. The solid lines show the same analysis when incorporating both the AT2017gfo $\mathcal{L}_{\rm{EM}}$ and GW170817 $\mathcal{L}_{\rm{GW}}$ data. Due to considerable flexibility in ejecta mass model systematics (i.e., $f_{\rm{disk}},\alpha_{\rm{dyn}}$), our model recovers once again a nearly indistinguishable distribution of ejecta parameters. Since our inference tries to fit AT2017gfo observations with our model, and since our ejecta model allows for systematic error, our inferences always find consistent solutions for the best-fitting ejecta, i.e. the ejecta parameters that consistently explain AT2017gfo within the context of our kilonova model family, subject to the limitations of those kilonova models. The much larger and more rapidly varying electromagnetic likelihood dominates our inference of these ejecta properties. By contrast, the GW information only slightly perturbs the overall conclusions obtained electromagnetically. The inclusion of an ejecta velocity model systematic parameter does not affect the similarity of the ejecta posteriors.
In Figure 4, we plot the light curves associated with the Nedora21 GW+EM posteriors from Figure 3, showing both the mean and posterior 68% credible interval. The width of these posterior intervals derives from and reflects the narrow posteriors shown in Figure 3 (including suppressed parameters reflecting model systematic uncertainty). As in previous studies (e.g. [72, 94]), we observe residual model systematics relative to the data, where the largest deviations between our models and the data occur in the bluer filters, namely the $g$ and $r$ bands, starting at $\sim 3$ days. Though we explore possible reasons for this blue filter discrepancy in prior work (see Figure 2 of [72]), we do not explicitly include this model fidelity issue in our formulation of $\sigma_{\rm{sys}}$ Both in this work, where we treat $\sigma_{\rm{sys}}$ as a free parameter, and in our previous EM-only inferences, where we assume $\sigma_{\rm{sys}}$ = 0.5 [72, 95, 52], the deviations that would arise from variations in the posteriors are well-encompassed by our systematics. Thus, for all intents and purposes, we expect the slight deviations of the posteriors in Figure 3 to have negligible effects on our findings.
@@
Figure 5: Binary parameter posterior distributions using EM (dashed) and GW+EM (solid) likelihoods using $\log f_{\rm{disk}}$ priors of [-0.9, -0.75], [-1.4, -1.2], and [-1.2, -1] for DiCo20, KruFo20, and Nedora21, respectively. The inclusion of the GW likelihood most noticeably affects the chirp mass $\mathcal{M}_{c}$ and mass asymmetry $\delta$, while simultaneously providing tighter constraints on the neutron star radius $R_{\rm{1.4}}$ and ejected disk fraction $f_{\rm{disk}}$
The corner plots in Figure 5 show 90% two-dimensional credible intervals and one-dimensional marginal posterior distributions for the underlying binary ($m_{c},\delta$), EOS ($R_{\rm{1.4}}$), and ejecta systematic ($f_{disk},\alpha_{dyn},\beta_{\phi}$) model parameters. As in Figure 3, the dashed curves represent inference using data from AT2017gfo alone and the solid curves consider both the AT2017gfo and GW170817 data. We do not plot the posteriors for $s_{\rm{1z}},s_{\rm{2z}}$, and $\sigma_{\rm{sys}}$ as they are largely uninteresting in the context of understanding the ejecta fits and would have reduced the legibility of Figure 5. Following a preliminary analysis to identify the posterior support for each ejecta fit, we perform focused analyses targeting narrow regions of $f_{\rm{disk}}$ for each ejecta fit: $\log f_{\rm{disk}}$ is uniform over [-0.9, -0.75], [-1.4, -1.2], and [-1.2, -1] for DiCo20, KruFo20, and Nedora21, respectively.
As expected, our analysis finds that the three ejecta models can only explain AT2017gfo using dramatically different progenitor binaries and EOS. Two of the three models prefer that the progenitor of AT2017gfo is a highly asymmetric binary ($\delta>0$) with smaller NS radius $R_{\rm{1.4}}$, with the DiCo20 models allowing equal-mass ($\delta=0$) binaries with exceptionally large NS radius. Conversely, these inferences suggest that a surprisingly small fraction ($\sim 4-20\%$) of the expected disk mass is ejected as a wind, in contrast to most previous work [96, 97], but consistent with predictions from three-dimensional general relativistic, full transport neutrino radiation magnetohydrodynamics disk simulations [98].
III.2 Allowing no uncertainty in ejecta velocity ($\beta_{\phi}=0$)
Figure 6: Same as Figures 2 and 5, but not allowing for systematic uncertainty in $v_{d}$ (i.e. $\beta_{\phi}=0$). Note the recreation of the $v_{\rm{ej}}$ contours from Figure 1 in the $R_{\rm{1.4}}$ vs. $\delta$ plane.

### v2

@@ II.3 The likelihood of ejecta parameters for AT2017gfo given electromagnetic observations
The AT2017gfo data is originally presented in [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 19, 84, 14, 85, 16, 17]. Each combination of ejecta parameters $\vec{x}_{\theta}$ is evaluated by the MLP to produce a light-curve prediction $\hat{y}$ for every one of the grizyJHK broadband filters. We calculate the residual between the MLP prediction $\hat{y}$ and the AT2017gfo observed data in time $i$ for every band $B$ by way of the reduced-$\chi^{2}$ statistic
@@ II.4 The likelihood of source parameters for GW710817 given gravitational wave observations
Following previous studies that constrain properties of GW170817 and the equation of state from GW observations [86, 87, 88, 89], we use the RIFT parameter inference engine [90, 91, 92] to explore and evaluate the marginal likelihood $L(x)$ for a wide range of different compact binary source parameters $x$. Following the RIFT paradigm, the parameters $x$ include all intrinsic (detector-frame) quantities needed to phenomenologically characterize the binary’s inspiral: its two component masses, spins, and tidal deformabilities, assumed independent of any specific EOS model. We then use standard nonparametric interpolation techniques (here, random forests from sklearn) to interpolate the marginal likelihood to provide a continuous estimate $\hat{L}(x)$. Our GW likelihood takes the form
[…]
with input parameters derived from $x~=~(m_{1},m_{2},\chi_{1},\chi_{2},\Lambda_{1},\Lambda_{2})$. To be more concrete, we will employ precisely the same analysis framework and settings as used in Ref. [93]. As a brief review to establish notation, gravitational wave observations of GW170817 constrain the binary masses $m_{i}$ and tidal deformabilities $\Lambda_{i}$ for each component $i$ of the binary [1, 86, 26, 94]. The presence of matter impacts the binary’s inspiral at leading order through each component’s dimensionless tidal deformability parameter $\Lambda_{i}=(2/3)k_{2}(c^{2}r_{i}/Gm_{i})^{5}$ [95], where $k_{2}$ is the $l=2$ Love number, $m_{i}$ is the neutron star mass, and $r_{i}$ is the radius of a $1.4M_{\odot}$ neutron star. The leading order contribution to the phase evolution of a GW inspiral is given by the weighted combination of $\Lambda_{i}$ terms
@@
The impacts of tidal effects on outgoing radiation are included in many conventional approximate phenomenological estimates of outgoing radiation from merging compact binaries. In this work, we begin with precisely the marginal likelihood data accumulated in Ref. [93]. That investigation was performed with a contemporary state-of-the-art model, IMRPhenomPv2_NRTidalv2 [96, 97], which incorporates precession physics but omits higher-order modes. This GW analysis was performed with open GW data for GW170817 available from GWOSC [98], using the same power spectral densities provided with GWTC-1 [99, 100], over a frequency range from $23{\rm Hz}$ to $1700{\rm Hz}$, with known sky location and source luminosity distance derived from the electromagnetically-identified host galaxy. For exploration purposes, most prior assumptions were customary (e.g., uniform in detector-frame component masses), noting that the previous investigation employed multiple spin priors and mass ratio priors to accumulate marginal likelihood information over a wide range of source parameters. Although this analysis included prior information about the alignment between the binary’s angular momentum direction and the line of sight, inferred from late-time radio afterglow observation [101, 102], this information should not significantly impact the GW intrinsic parameters (the binary component masses, spins, and tidal deformabilities).
Because the original marginal likelihood grid accumulated in Ref. [93] did not extend to cover all of the extreme BNS configurations needed for this study, we replicated and extended this analysis to ensure that marginal likelihood data was present whenever required for joint EM/GW analysis. For simplicity, we employ the same GW model throughout for all $\tilde{\Lambda}$, keeping in mind that this model was not calibrated over the full range of $\Lambda$ needed to carry out our investigation.
@@ II.5 Bayesian inference
We use the RIFT framework [103] to adaptively perform Monte Carlo integration and generate samples using a joint likelihood $\ln\mathcal{L}=\ln\mathcal{L}_{\rm{GW}}+\ln\mathcal{L}_{\rm{EM}}$. In summing the separate GW and EM likelihoods, we assume their equal contribution to the joint likelihood and that our error model is well understood and lacking additional systematics (both modeling and data). While possible choices for composite likelihood weights have been discussed in the literature [104], in this work we stick to established conventions and leave the exploration of weighted composite likelihoods to a separate study. As previously, we employ an adaptive volume Monte Carlo integrator, following closely the approach outlined in Ref. [105]. The adaptive volume integrator allows for more efficient sampling given the higher-dimensional space explored in this work. As our inference is performed via adaptive Monte Carlo integration, the reliability of our posterior can be expressed in terms of a number of effective samples $n_{\rm eff}$. Several different conventions exist for this number; see the appendix of Ref. [92] for more detailed discussion. For this study, we terminate our analyses when $n_{\rm eff}\simeq 10^{3}$, indicating sufficient convergence.
Table 1 shows the sampling parameters used during inference, along with their associated lower and upper limits and priors. The combination of the chirp mass ${\cal M}_{c}$ and mass asymmetry $\delta$ priors provide an equivalent locally uniform prior for the binary masses $m_{1}$ and $m_{2}$. In addition to the masses, we also sample for the spins of the neutron stars $s_{\rm{1z}}$ and $s_{\rm{2z}}$. We assume that both of the neutron stars have the same radius given a 1.4 $M_{\odot}$ neutron star $R_{1.4}$, treating this radius as a proxy for the nuclear EOS. As described in Sec. II.2, we introduce parameters $\alpha_{\rm{dyn}}$, $f_{\rm{disk}}$, and $\beta_{\phi}$ that encapsulate the uncertainty in our dynamical ejecta mass, wind ejecta mass, and dynamical ejecta velocity, respectively. These parameters are sampled in addition to the light-curve systematics $\sigma_{\rm{sys}}$, effectively allowing for three separate measures of systematic uncertainty in our analysis. Finally, we sample for the viewing angle $\theta$ as in prior work (e.g. Ref. [56]). All parameters in Table 1 are assumed to be dimensionless unless their names are followed by units in brackets.
@@
$\mathcal{M}_{c}$ $M_{\odot}$] [1.1853, 1.1880] Jointly uniform $m_{1},m_{2}$
$\delta$ [0,0.8]{[0,0.8]} Jointly uniform $m_{1},m_{2}$
@@
$R_{1.4}$ [km] [$8,50$] Uniform
$\log\alpha_{\rm{dyn}}$ [5,5-5,5] Uniform
$\log f_{\rm{disk}}$ [5,0-5,0] Uniform
$\beta_{\phi}$ [2,3-2,3] Uniform
@@ III Results
For all the analyses in this section, we assume that the dynamical ejecta mass systematic uncertainty $\alpha_{\rm{dyn}}$ and light-curve systematic uncertainty $\sigma_{\rm{sys}}$ are treated as free parameters. In Section III.1, we treat $f_{\rm{disk}}$ and $\beta_{\phi}$ as free parameters to characterize the uncertainties associated with the wind ejecta mass (derived from the ejected disk mass) and ejecta velocity, respectively. In Section III.2, we fix $\beta_{\phi}=0$ to show the effect of ignoring ejecta velocity systematics during inference. Finally, in Section III.3, we fix both $\beta_{\phi}=0$ and $f_{\rm{disk}}=0.30$ to highlight how making assumptions about no ejecta velocity or disk ejection fraction systematics renders meaningful interpretation of our inference results intractable.
III.1 All systematics included ($\alpha_{\rm{dyn}},f_{\rm{disk}},\beta_{\phi}$ free)
Figure 3: Binary parameter posterior distributions using EM (dashed) and GW+EM (solid) likelihoods using $\log f_{\rm{disk}}$ priors of [-0.9, -0.75], [-1.4, -1.2], and [-1.2, -1] for DiCo20, KruFo20, and Nedora21, respectively. The inclusion of the GW likelihood most noticeably affects the chirp mass $\mathcal{M}_{c}$ and mass asymmetry $\delta$, while simultaneously providing tighter constraints on the neutron star radius $R_{\rm{1.4}}$ and ejected disk fraction $f_{\rm{disk}}$. Gray shaded regions represent $\delta>0.3$ for which the secondary star’s mass is sub-solar ($1M_{\odot}$).
The corner plots in Figure 3 show 90% two-dimensional credible intervals and one-dimensional marginal posterior distributions for the underlying binary ($\mathcal{M}_{c},\delta$), EOS ($R_{\rm{1.4}}$), and ejecta systematic ($f_{disk},\alpha_{dyn},\beta_{\phi}$) model parameters. The dashed black, red, and blue lines correspond to inferences using only AT2017gfo data $\mathcal{L}_{\rm{EM}}$ for the KruFo20, DiCo20, and Nedora21 ejecta fits, respectively. The solid lines show the same analysis when incorporating both the AT2017gfo $\mathcal{L}_{\rm{EM}}$ and GW170817 $\mathcal{L}_{\rm{GW}}$ data. We do not plot the posteriors for $s_{\rm{1z}},s_{\rm{2z}}$, and $\sigma_{\rm{sys}}$ as they are largely uninteresting in the context of understanding the ejecta fits and would have reduced the legibility of the figure. Following a preliminary analysis to identify the posterior support for each ejecta fit, we perform focused analyses targeting narrow regions of $f_{\rm{disk}}$ for each ejecta fit: $\log f_{\rm{disk}}$ is uniform over [-0.9, -0.75], [-1.4, -1.2], and [-1.2, -1] for DiCo20, KruFo20, and Nedora21, respectively.
We verify that our ejecta fit posteriors produce consistent kilonova ejecta properties in Figure 4. We begin by examining the posteriors for the binary’s ejecta properties, primarily constrained by our EM likelihood $\mathcal{L}_{\rm{EM}}$. Figure 4 shows the predicted ejected masses, bulk velocity (attributed to the dynamical ejecta), and the viewing angle derived from the inferences presented in Figure 3. In this analysis, we adopt $v_{w}=0.10c$, consistent with the viscosity-driven wind ejecta in numerical simulations and our previous analysis of AT2017gfo [56]. As in Figure 3, the dashed curves represent inference using data from AT2017gfo alone and the solid curves consider both the AT2017gfo and GW170817 data. Due to considerable flexibility in ejecta mass model systematics (i.e., $f_{\rm{disk}},\alpha_{\rm{dyn}},\beta_{\phi}$), our model recovers once again a nearly indistinguishable distribution of ejecta parameters. Since our inference tries to fit AT2017gfo observations with our model, and since our ejecta model allows for systematic error, our inferences always find consistent solutions for the best-fitting ejecta, i.e. the ejecta parameters that consistently explain AT2017gfo within the context of our kilonova model family, subject to the limitations of those kilonova models. The much larger and more rapidly varying electromagnetic likelihood dominates our inference of these ejecta properties. By contrast, the GW information only slightly perturbs the overall conclusions obtained electromagnetically. The inclusion of an ejecta velocity systematic parameter does not affect the similarity of the ejecta posteriors. As in Ref. [56], our $\theta$ posteriors deviate appreciably from the prior, owing to our model reconciling the tension between the “blue” and “red” emission by sampling away from the $\theta$ prior (see their Fig. 10).
In Figure 5, we plot the light curves associated with the Nedora21 GW+EM posteriors from Figure 4, showing both the mean and posterior 68% credible interval. The width of these posterior intervals derives from and reflects the narrow posteriors shown in Figure 4 (including suppressed parameters reflecting model systematic uncertainty). As in previous studies (e.g. [77, 106]), we observe residual model systematics relative to the data, where the largest deviations between our models and the data occur in the bluer filters, namely the $g$ and $r$ bands, starting at $\sim 3$ days. Though we explore possible reasons for this blue filter discrepancy in prior work (see Figure 2 of [77]), we do not explicitly include this model fidelity issue in our formulation of $\sigma_{\rm{sys}}$. We note that a better fit to late-time blue photometry would most likely invoke a slightly more massive, slower wind ejecta component; a slower wind would produce a longer-lived blue kilonova, with the increased mass accounting for the flux loss from slower-moving ejecta (assuming ejecta under local thermodynamic equilibrium (LTE), see Ref. [107] for non-LTE effects on late-time emission). Both in this work, where we treat $\sigma_{\rm{sys}}$ as a free parameter, and in our previous EM-only inferences, where we assume $\sigma_{\rm{sys}}$ = 0.5 [77, 108, 56], the deviations that would arise from variations in the posteriors are well-encompassed by our systematics.
@@ III.1 All systematics included ($\alpha_{\rm{dyn}},f_{\rm{disk}},\beta_{\phi}$ free)
As expected, our analysis finds that the three ejecta models can only explain AT2017gfo using dramatically different progenitor binaries and EOS. Two of the three models prefer that the progenitor of AT2017gfo is a highly asymmetric binary ($\delta>0.2\implies q<0.66$) with smaller NS radius $R_{\rm{1.4}}$, with the DiCo20 models allowing equal-mass ($\delta=0$) binaries with exceptionally large NS radius. Conversely, these inferences suggest that a surprisingly small fraction ($\sim 4-20\%$) of the expected disk mass is ejected as a wind, in contrast to most previous work [109, 110], but consistent with predictions from three-dimensional general relativistic magnetohydrodynamic (GRMHD) disk simulations with full transport neutrino radiation [111] (although past 1 second, these simulations have a larger fraction of ejected mass [112]; more late-time ($t>1$ sec) disk simulations are required for a conclusive comparison).
III.2 Allowing no uncertainty in ejecta velocity ($\alpha_{\rm{dyn}},f_{\rm{disk}}$ free, $\beta_{\phi}=0$)
Figure 6: Same as Figures 2 and 3, but not allowing for systematic uncertainty in $v_{d}$ (i.e. $\beta_{\phi}=0$). Note the recreation of the $v_{\rm{ej}}$ contours from Figure 1 in the $R_{\rm{1.4}}$ vs. $\delta$ plane.